OpenStudy (anonymous):

What is the solution to the equation log2 x + log2 4 = 3 ?

OpenStudy (anonymous):

x = 2 x = 128 x = 32 x = 1/2

OpenStudy (helder_edwin):

if you equation \[ \LARGE \log^2x+\log^24=3 \] or \[ \LARGE \log_2x+\log_24=3 \]

OpenStudy (anonymous):

it's the second @helder_edwin . Answer is x = 2

OpenStudy (anonymous):

the bottom one

OpenStudy (anonymous):

how did you get tht?

OpenStudy (helder_edwin):

OK do you want the work or the answer. @Diogo gave you the answer

OpenStudy (anonymous):

i want the work

OpenStudy (helder_edwin):

OK you have \[ \LARGE 3=\log_2x+\log_24=\log_2x+\log_22^2=\log_2x+2 \] do u get it?

OpenStudy (anonymous):

put everything in log10 form. example: log(x)/log(2) + 2 = 3 log(x) = log(2) x = 2

OpenStudy (anonymous):

ohh okay i got it ! you guys explain it better then my book

OpenStudy (campbell_st):

\[\log _{2}4 = \log _{2} 2^2 = 2\] so the equation is \[\log _{2} x + 2 = 3... or ... \log _{2 } x = 1\] using the log law \[\log _{a}b = c ........ a^{\log _{a} b} = a^c ... or.... b = a^c\] then \[x = 2^1\] or x = 2